Systems Staff Meeting Notes
May 16, 2008
May 16, 2008
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Chris Nicolich, Amy Polk, Stephen WilliamsAnnouncements
Ongoing Systems Updates
- Newberry Watermelon festival is this weekend.
- "Indiana Jones" comes out in one week.
- "Sex & the City" comes out in two weeks.
- Support Services Recap: floors in West - supplies did not arrive. Will probably replace it during the summer break before fall semester. We are waiting to schedule the work until we have the supplies in hand this time.
Maps is relocating due to work on the MSL plaza area - crack in the concrete a source of perpetual leaks. They may need to store public machines in the Systems vault.
AFA power - Rob has the key. He is going to get the power checked.
A fire drill is scheduled for next week. Remember to meet at the Murphree statue.
Library Awards Program has been announced. Don't forget to nominate your colleagues.
- Orders - IBM/Maps purchase - we added one more hard drive to the order.
- MSL laptops will ship soon. We are adding RAM - that will also be shipping soon.
- Training next week. Eight people are registered. Tues-Thurs 1-5PM Room 419 East.
- Systems training - 419 West - Stephen or Chris will reserve it.
Software Projects Status
- Cynthia is trying to network a printer on NT25. Logan suggests putting it on the virtual server.
- Smathers2K3DC1 - was having replication problems. Logan took it down. When it was brought back up, it did not promote correctly. It was removed from AD and Logan made a new SL2K3DC1 to replace it. No errors, so we hope the problem is solved. Will promote a WINS server on it. Logan re-did the WINS server set-up. He will put one in each location: East, West, MSL, Education, AFA, etc.
- SX280 issue - Dell management is reviewing whether they will replace the motherboards that are under warranty and ship motherboards for those that are out-of-warranty.
- Cynthia mentions she has gotten 4 tickets citing problem motherboards - only one was actually a problem. Send the three functional ones back to the users. Will wants to look at the one with the valid problem.
- NT2R outage Wednesday - cluster name would not come online. Will created a new one.
- McAfee update - going pretty well. Couple machines are not cooperating.
- CJK - machine not delivered to Logan yet.
- Ticket from Laurie Brennan this morning, reporting that the servers went off for a few minutes in MSL. They are keeping a log, as this has happened before.
DLC Greenstone Issues
- Wikis - nothing new. Matthew logged in a ticket to Logan to set one up.
- Grover - will reserve a room for training next Thursday.
- Affiliation Modification Proposal - Winston is working on APIs - testing access.
Students and Departments
- Greenstone - problem - it is totally down. Seems to be related to domain controller issues. Logan to check on the DNS.
- Greenstone backups - will fail. When it is still trying to run, it will back up the rest of the jobs. QuickErase is recommended.
Systems Security Management
- Rob will be working during summer semester.
- Meeting held.